# Design and Performance Comparison of FuzzyPID and Non-linear Model Predictive Controller for 4-Wheel Omni-drive Robot

### v1

@@ Abstract
Trajectory tracking for an Omni-drive robot presents a challenging task that demands an efficient controller design. To address the limitations of manual tuning, we introduce a self-optimizing controller named fuzzyPID, leveraging the analysis of responses from various dynamic and static systems. The rule-based controller design is implemented using Matlab/Simulink, and trajectory tracking simulations are conducted within the CoppeliaSim environment. Similarly, a non-linear model predictive controller(NMPC) is proposed to compare tracking performance with fuzzyPID. We also assess the impact of tunable parameters of NMPC on its tracking accuracy. Simulation results validate the precision and effectiveness of NMPC over fuzzyPID controller while trading computational complexity.
@@
fuzzyPID controller, Omni drive, trajectory tracking, kinematics; NMPC, Center of Gravity(COG)
@@ 1 Introduction
Conventional PID controllers have been widely employed in mobile robotics for efficient speed control [5] and waypoint tracking [6]. However, optimizing these controllers for desirable output characteristics can be challenging, and they may not be suitable for a wide range of operating conditions [5, 7]. Consequently, an alternative and effective approach to controller design is utilizing fuzzy logic. Fuzzy logic controllers leverage sets of rules derived from the responses of various static and dynamic systems to adjust tuning parameters based on fuzzy input variables [8, 9]. These controllers exhibit good system performance, transient response, and disturbance rejection capabilities. Moreover, they have been successfully applied to various non-linear plants that are challenging to model due to a lack of sufficient parameters [10, 11]. From a driving perspective, studies such as [12, 13] have successfully evaluated the performance of path planners designed based on fuzzy logic for Omni-drive robots.
@@
With the advancement of powerful computing devices, optimization-based control techniques have become integral in various autonomous mobile industries [14]. Model Predictive Control (MPC) is one such method widely used for path tracking of mobile robots, as it takes into consideration various constraints for optimal control inputs [15, 16, 17, 18, 19]. In comparison with traditional PID methods, MPC offers higher accuracy, smoother control inputs, and increased resistance to external disturbances. Papers such as [20, 16] also highlight the superior performance of MPC over PID controllers, especially when considering kinematic constraints in mobile robots. Considering this, different variants of MPC-based path tracking controllers have been evaluated in recent years, based on the nature of the prediction model’s nonlinearity and the representation of output state quantities [21, 22]. In 2007, the paper by Conceiccao et al. [23] proposed non-linear MPC based on error kinematics models for Omni-drive robots for path following, which was particularly challenging due to hardware constraints at that time. To address these challenges, linear MPC was introduced, which linearized the robot model to simplify computation operations. In a subsequent paper by Wang et al. [24], the authors designed linear MPC and demonstrated its effectiveness for point stabilization and trajectory tracking. Similarly, Kanjanawanishkul et al. [17] presented and applied linear error MPC in real-time for Omni-drive robots, considering system and input constraints to trade for stability. Furthermore, Grüne et al. [25] introduced various extensions of NMPC for time-varying references with or without terminal constraints.
We utilize CoppeliaSim as the simulation platform to conduct our experiments. We design the physical model of a four-wheel omni-drive robot in SolidWorks and import the URDF format into CoppeliaSim. Subsequently, it is interfaced with Matlab/Simulink via an external API to enable control. Physical parameters such as frictional force and air resistance are maintained at default values in the simulation environment.
@@
In summary, our paper comprises three main sections. The initial two sections are dedicated to modeling the kinematics of the robot and subsequently addressing path planning. Following this, we introduce an Omni-drive controller, employing two distinct methods: fuzzy logic with the Mamdani model and Non-linear Model Predictive Control (NMPC). We proceed to assess and compare the performance of these approaches in guiding the robot along the reference trajectory generated by the path planning algorithms. The design of the Omni-drive controller involves formulating a non-linear predictive model, treated as a cost minimization problem. The objective is to determine optimal control inputs that adhere to bounded input constraints, subsequently applied to the robot kinematics. This ensures the robot effectively follows the desired path outlined by the path planning algorithms.
@@ 2 Kinematics model
Figure 1: Component division of translational velocity of omni wheel
@@ 3 Path Planning
We select the $A^{*}$ algorithm for its efficiency and its appropriateness in devising optimal paths for static environments [3]. This algorithm uses the heuristic approach to estimate the best optimal path by avoiding obstacles in the given 2D environment if the solution really exists. Here, the environment of the robot is discretized into a number of small grid cells called an occupancy grid map filled with the binary digit of ’1’ and ’0’ which denotes the obstacle and free space respectively. In the following paper [27], different heuristic functions were compared that best optimize the algorithm. But, the simple way is to use Manhattan distance which is based upon the sum of the absolute differences between the two vectors i.e current node to goal node which is given as,
@@ 4.1 FuzzyPID controller
Figure 4: The range of triangular membership function are adjusted based on input and output variables. Both input variables $e$ and $de$ are in the range between -3 to 3. For output variables ($kp,ki,kd$), range is adjusted in between -0.3 to 0.3 keeping the structure of membership function similar.
@@
The design of the fuzzy Controller begins with defining input variables $e(k)$ and the change in error $de(k)$. Initially, these are considered as crisp inputs and need to be converted into 7 overlapping fuzzy sets through the process of fuzzification. The universe of discourse for the variables is defined based on the ranges of the input variables. Triangular membership functions are employed for the fuzzification step, where the membership values (degree of truth) are obtained in the range of 0 to 1 on the universe of discourse for the given input values. This process is illustrated in Fig. 4. Fuzzy sets are represented in linguistic form using the set of variables NB, NM, NS, ZO, PS, PM, and PB, where N, Z, H represent Negative, Zero, and Positive respectively, and B, M, S, O represent Big, Medium, Small, and Zero respectively. The next step involves designing the Mamdani-based fuzzy inference system, where a set of fuzzy rules is defined based on the relationship between fuzzy input sets ($e(k)$ and $de(k)$) and output parameters ($kp,ki,kd$). These rules are formulated by control engineers based on their experience with various system responses. In this context, we establish 49 rules for each fuzzyPID output parameter, drawing reference from the work in [28]. A fuzzy rule takes the form of an IF(antecedent)-THEN(consequent) statement, such as ”if $e(k)$ is PB and $de(k)$ is PB, then apply a large negative $kp$ (i.e., NB)”. The process is executed by fuzzy inference engine that employ the max-min composition.
@@
The final step in the design of the fuzzyPID controller is defuzzification process, which involves the conversion of linguistic variables into crisp output values using membership functions. This step is the reverse process of fuzzification, where values between 0 and 1 are converted using the Center of Gravity (COG) defuzzification technique. Ultimately, we obtain changes in the PID parameters, which are added to the previous PID parameters to continuously track the reference values, expressed as $K_{pid}=K_{pid}^{\prime}+\Delta K_{pid}$. The entire fuzzyPID controller is implemented in Simulink, as illustrated in Fig. 5.
Figure 5: Simulink model of fuzzyPID
@@ 4.2 Non-linear MPC
Figure 6: Comparison of tracking performance of pose(XY) for both fuzzyPID and NMPC with prediction horizon of 15 with a total tracking time of 30 seconds
Figure 6: Comparison of tracking performance of pose(XY) for both fuzzyPID and NMPC with prediction horizon of 15 with a total tracking time of 30 seconds
Figure 7: Comparison of tracking performance of pose($\theta$) for both fuzzyPID and NMPC with prediction horizon of 15 with a total tracking time of 30 seconds
@@
$\left\{\begin{array}[]{c}\mathbf{e}(k\mid k)=\mathbf{\bar{x}}(k\mid k)-\mathbf{x}_{ref}(k\mid k)\\ \vdots\\ \mathbf{e}\left(k+N_{p}\mid k\right)=\mathbf{\bar{x}}\left(k+N_{p}\mid t\right)-\mathbf{x}_{ref}\left(k+N_{p}\mid k\right)\end{array}\right.$ (24)
@@ 5 Simulation setup
We conduct trajectory tracking experiments using CoppeliaSim, with our primary objective being the comparison of tracking accuracy between both controllers while keeping the sampling time constant. The tracking performance is simulated over a fixed time intervals of 20 and 30 seconds for the generated trajectory, and the results obtained are plotted for a fair comparison. The reference and actual 2D-pose values for both the controllers are shown in Fig. 7 and 7. For NMPC, we use a prediction horizon value of 15. It is crucial to note that altering the prediction horizon value has adverse effects on the controller’s efficiency, as illustrated in Fig. 9 and 9. The accuracy of the controller is also dependent on the number of intermediate poses sampled within the given tracking time. In this trajectory experiment, we choose the parameters as, $\mathbf{Q}=15\mathbb{I}_{3}$, $\mathbf{R}=\mathbb{I}_{2}$, $t_{s}=0.1s$, $v_{max}=1.5m/s$, $\omega_{max}=3.14rad/s$.
We do not take hardware constraints into account as limitations for our simulation platforms. Mean Cross-track Error (ME) and Mean Absolute Error (MAE) metrics serve as the evaluation criteria for judging our results. ME is calculated by averaging the Euclidean distance between the target pose and the real pose at each sampling period. Similarly, we calculate the average value of all absolute deviations between the reference heading angle and the robot heading angle for MAE. Lower error values correspond to better tracking accuracy. Furthermore, we compare the tracking results for different prediction horizon values ($N_{p}$) in the case of Non-linear Model Predictive Control.
Table 1: Comparison of tracking error of omni drive robot for both fuzzyPID and NMPC
fuzzyPID NPMC($N_{p}=15$)
Tracking time ME($XY$) MAE($\theta$) ME($XY$) MAE($\theta$)
20 seconds 0.0712 m 0.0793 rad 0.0529 m 0.0564 rad
30 seconds 0.0499 m 0.0613 rad 0.0439 m 0.0391 rad
6 Comparison of fuzzyPID and NMPC
Table 1 presents the tracking error for both fuzzyPID and NMPC with a prediction horizon value of 15. The NMPC controller achieves a lower error compared to fuzzyPID in both metrics. As the tracking time increases, a corresponding improvement in tracking accuracy is observed for both controllers. FuzzyPID is identified as more computationally efficient than NMPC, albeit at the expense of tracking accuracy.

### v2

# Comparative Analysis of NMPC and Fuzzy PID Controllers for Trajectory Tracking in Omni-Drive Robots: Design, Simulation, and Performance Evaluation

@@ Abstract
Trajectory tracking for an Omni-drive robot presents a challenging task that demands an efficient controller design. This paper introduces a self-optimizing controller, Type-1 fuzzyPID, which leverages dynamic and static system response analysis to overcome the limitations of manual tuning. To account for system uncertainties, an Interval Type-2 fuzzyPID controller is also developed. Both controllers are designed using Matlab/Simulink and tested through trajectory tracking simulations in the CoppeliaSim environment. Additionally, a non-linear model predictive controller (NMPC) is proposed and compared against the fuzzyPID controllers. The impact of tunable parameters on NMPC’s tracking accuracy is thoroughly examined. We also present plots of the step-response characteristics and noise rejection experiments for each controller. Simulation results validate the precision and effectiveness of NMPC over fuzzyPID controllers while trading computational complexity. Access to code and simulation environment is available in the following link: https://github.com/love481/Omni-drive-robot-Simulation.git.
@@
Omni-drive robot, fuzzyPID controller, NMPC, Center of Gravity(COG), Path-tracking, Step-response, Noise rejection
@@ 1 Introduction
Conventional PID controllers [5] have been widely employed in various robotics application for efficient speed control [6] and waypoint tracking [7]. However, optimizing these controllers for desirable output characteristics can be challenging, and they may not be suitable for a wide range of operating conditions [6, 8]. Consequently, an alternative and effective approach to controller design is utilizing fuzzy logic. Fuzzy logic controllers leverage sets of rules derived from the responses of various static and dynamic systems to adjust tuning parameters based on fuzzy input variables [9, 10, 11, 12]. These controllers exhibit good system performance, transient response, and disturbance rejection capabilities. Moreover, they have been successfully applied to various non-linear plants that are challenging to model due to a lack of sufficient parameters [13, 14].
In the context of autonomous path tracking, a fuzzy-based controller has been proposed that leverages the geometric properties of the path ahead for a differential drive robot [15]. This controller considers the curvature and distance to the next bend relative to the robot’s current location, thereby mimicking real driving behavior to determine the desired cruise speed. Similarly, T1-Fuzzy PID controllers have been developed using 25 fuzzy rules for the automatic tuning of PID gains in differential drive robots and quadcopters [7, 16]. The tuned gains are then used to compute the desired velocity, guiding the robot along the reference trajectory and minimizing tracking errors, which demonstrated superior performance compared to classical PID controllers. A Particle Swarm Optimization (PSO)-optimized fuzzy controller has been introduced that directly outputs joint torques for a 2-DOF planar robot [17]. In this approach, PSO is employed to optimize the antecedent and consequent parameters of the Mamdani-based fuzzy logic system. For omni-drive robots, studies have successfully evaluated the performance of fuzzy logic-based path planners [18, 19, 20]. These approaches continuously update control signals derived from the fuzzy logic controller based on deviation errors, guiding the robot to the desired waypoints.
Recently, significant research has focused on developing effective Type-2 Fuzzy Logic Controllers (FLCs) to model various sources of uncertainties in nonlinear systems [12, 21, 22]. A Type-2 Fuzzy Logic Controller has been presented for modeling uncertainties in measurements during the tracking of mobile objects in robotic soccer games [23]. This controller uses angle error and change in angle as inputs to the fuzzy system, outputting speeds and directions to control the motion of the agents. Similarly, a novel application of GA-optimized IT2-FPD+I controllers has been explored for 5-DOF redundant robot manipulators to track desired joint trajectories [24]. Their experiments demonstrated robustness in terms of disturbance rejection, model uncertainties, and noise injections, outperforming previous T1-FPD and conventional PID controllers. A new method for building an optimal structure of a PID-type IT2 FLC system has been proposed for controlling joint actions in delta robots [25]. This approach uses the Non-dominated Sorting Genetic Algorithm (NSGA) to tune the controller’s scaling factors by formulating the problem as a multi-objective optimization task. In another study, Social Spider Optimization (SSO) has been utilized for parameter tuning [26]. Additionally, the effects of Footprint of Uncertainty (FOU) parameters on generating the control surface for a Single Input IT2-Fuzzy PID (SI-IT2-FPID) controller have been explained, with performance demonstrations for trajectory tracking of UAVs [27].
@@
With the advancement of powerful computing devices, optimization-based control techniques have become integral in various autonomous mobile industries [28]. Model Predictive Control (MPC) is one such method widely used for path tracking of mobile robots, as it takes into consideration various constraints for optimal control inputs [29, 30, 31, 32, 33]. In comparison with traditional PID methods, MPC offers higher accuracy, smoother control inputs, and increased resistance to external disturbances. Papers such as [34, 30] also highlight the superior performance of MPC over PID controllers, especially when considering kinematic constraints in mobile robots. Considering this, different variants of MPC-based path tracking controllers have been evaluated in recent years, based on the nature of the prediction model’s nonlinearity and the representation of output state quantities [35, 36]. In 2007, the paper by Conceiccao et al. [37] proposed non-linear MPC based on error kinematics models for Omni-drive robots for path following, which was particularly challenging due to hardware constraints at that time. To address these challenges, linear MPC was introduced, which linearized the robot model to simplify computation operations. In a subsequent paper by Wang et al. [38], the authors designed linear MPC and demonstrated its effectiveness for point stabilization and trajectory tracking. Similarly, Kanjanawanishkul et al. [31] presented and applied linear error MPC in real-time for Omni-drive robots, considering system and input constraints to trade for stability. Furthermore, Grüne et al. [39] introduced various extensions of NMPC for time-varying references with or without terminal constraints. Despite advancements in trajectory tracking, few approaches have been applied to Omni-directional robots, leaving a gap in proposing Nonlinear Model Predictive Control (NMPC) and assessing its performance against other control methods.
@@
We utilize CoppeliaSim as the simulation platform to conduct our experiments. We design the physical model of a four-wheel omni-drive robot in SolidWorks and import the URDF format into CoppeliaSim. Subsequently, it is interfaced with Matlab/Simulink via an external API to enable control. Physical parameters such as frictional force and air resistance are maintained at default values in the simulation environment. In summary, our paper comprises three main sections. The initial two sections are dedicated to modeling the kinematics of the robot and subsequently addressing path planning. Following this, we introduce an Omni-drive controller, employing two distinct methods: fuzzy logic with the Mamdani model and Non-linear Model Predictive Control (NMPC). We proceed to assess and compare the performance of these approaches in guiding the robot along the reference trajectory generated by the path planning algorithms. The design of the Omni-drive controller involves formulating a non-linear predictive model, treated as a cost minimization problem. The objective is to determine optimal control inputs that adhere to bounded input constraints, subsequently applied to the robot kinematics. This ensures the robot effectively follows the desired path outlined by the path planning algorithms.
@@ 3 Path Planning
We select the $A^{*}$ algorithm for its efficiency and its appropriateness in devising optimal paths for static environments [3]. This algorithm uses the heuristic approach to estimate the best optimal path by avoiding obstacles in the given 2D environment if the solution really exists. Here, the environment of the robot is discretized into a number of small grid cells called an occupancy grid map filled with the binary digit of ’1’ and ’0’ which denotes the obstacle and free space respectively. In Coppeliasim, we have created the custom static obstacles and build the map by treating robot as a point object. So, this increases the map area around the obstacles thereby easing the path planning problem for $A^{*}$ algorithm. In the following paper [41], different heuristic functions were compared that best optimize the algorithm. But, the simple way is to use Manhattan distance which is based upon the sum of the absolute differences between the two vectors i.e current node to goal node which is given as,
@@ 4.1 FuzzyPID controller
Figure 4: The range of triangular membership function are adjusted based on input and output variables. Both input variables $e$ and $de$ are in the range between -1 to 1. For output variables ($kp,ki,kd$), range is adjusted in between -0.1 to 0.1 keeping the structure of membership function similar.
@@
The design of the type-1 fuzzy Controller begins with defining input variables $e(k)$ and the change in error $de(k)$. Initially, these are considered as crisp inputs and need to be converted into 7 overlapping fuzzy sets through the process of fuzzification. The universe of discourse for the variables is defined based on the ranges of the input variables. Triangular membership functions are employed for the fuzzification step, where the membership values (degree of truth) are obtained in the range of 0 to 1 on the universe of discourse for the given input values. This process is illustrated in Fig. 4. Fuzzy sets are represented in linguistic form using the set of variables NB, NM, NS, ZO, PS, PM, and PB, where N, Z, H represent Negative, Zero, and Positive respectively, and B, M, S, O represent Big, Medium, Small, and Zero respectively.
The next step involves designing the Mamdani-based fuzzy inference system, where a set of fuzzy rules is defined based on the relationship between fuzzy input sets ($e(k)$ and $de(k)$) and output parameters ($kp,ki,kd$). These rules are formulated by control engineers based on their experience with various system responses. In this context, we establish 49 rules for each T1-fuzzyPID output parameter, drawing reference from the work in [42]. A fuzzy rule takes the form of an IF(antecedent)-THEN(consequent) statement, such as ”if $e(k)$ is PB and $de(k)$ is PB, then apply a large negative $kp$ (i.e., NB)”. The process is executed by fuzzy inference engine that employ the max-min composition. The fuzzy rules for each parameter are presented in Table 1.
[…]
@@
The final step in the design of the type-1 fuzzyPID controller is defuzzification process, which involves the conversion of linguistic variables into crisp output values using membership functions. This step is the reverse process of fuzzification, where values between 0 and 1 are converted using the Center of Gravity (COG) defuzzification technique. Ultimately, we obtain changes in the PID parameters, which are added to the previous PID parameters to continuously track the reference values, expressed as $K_{pid}=K_{pid}^{\prime}+\Delta K_{pid}$.
We designed the IT2-Fuzzy PID controller for trajectory tracking in an omnidirectional robot by applying the concepts outlined in Mendel’s work on uncertain systems [12]. The key difference between our IT2-FPID and previous controllers lies in the use of Interval Type-2 Fuzzy Sets (IT2 FS), along with type-reduction methods that convert IT2 FS to Type-1 Fuzzy Sets (T1 FS). In this design, IT2 FS are used to represent the fuzzy input and output variables, where uncertainties are captured by the Footprint of Uncertainty (FOU), providing an additional degree of freedom. For our implementation, the error $e(k)$ and its derivative $de(k)$ are represented by seven overlapping FOUs, as illustrated in Fig. 4. The upper membership functions (UMFs) of these FOUs are identical to the T1 FS, while the scaling factor for the lower membership functions (LMFs) and the lag value are set to 1.0 and 0.3, respectively. Additionally, we adopted the same rule base used in the T1 FPID controller. The firing intervals for each rule are calculated using the minimum implication method. The Karnik-Mendel (KM) algorithms are then employed to obtain T1 FSs by calculating the centroids using the iterative center of sets (COS) type-reduction method. Finally, defuzzification is performed by averaging the left and right points of the type-reduced set to produce the crisp output(PID gains).
@@ 4.2 Non-linear MPC
Figure 5: Comparison of tracking performance of pose(XY) for both type fuzzyPIDs and NMPC with prediction horizon of 15 with a total tracking time of 30 seconds
Figure 6: Comparison of tracking performance of pose($\theta$) for both type fuzzyPIDs and NMPC with prediction horizon of 15 with a total tracking time of 30 seconds
@@
$\left\{\begin{array}[]{c}\mathbf{e}(k\mid k)=\mathbf{\bar{x}}(k\mid k)-\mathbf{x}_{ref}(k\mid k)\\ \vdots\\ \mathbf{e}\left(k+N_{p}\mid k\right)=\mathbf{\bar{x}}\left(k+N_{p}\mid k\right)-\mathbf{x}_{ref}\left(k+N_{p}\mid k\right)\end{array}\right.$ (13)
@@ 5 Simulation setup
We conduct trajectory tracking experiments using CoppeliaSim, with our primary objective being the comparison of tracking accuracy between three controllers while keeping the sampling time constant. The tracking performance is simulated over a fixed time intervals of 20 and 30 seconds for the generated trajectory, and the results obtained are plotted for a fair comparison. We also assess the robustness of our proposed controllers in the presence of external noise. The reference and actual 2D-pose values for all the controllers are shown in Fig. 6 and 6. For NMPC, we use a prediction horizon value of 15. It is crucial to note that altering the prediction horizon value has adverse effects on the controller’s efficiency, as illustrated in Fig. 8 and 8. The accuracy of the controller is also dependent on the number of intermediate poses sampled within the given tracking time. In this trajectory experiment, we choose the parameters as, $\mathbf{Q}=15\mathbb{I}_{3}$, $\mathbf{R}=\mathbb{I}_{2}$, $t_{s}=0.1s$, $v_{max}=1.5m/s$, $\omega_{max}=3.14rad/s$.
Figure 9: Step response curve of our controllers
We do not take hardware constraints into account as limitations for our simulation platforms. Mean Cross-track Error (ME) and Mean Absolute Error (MAE) metrics serve as the evaluation criteria for judging our results. ME is calculated by averaging the Euclidean distance between the target pose and the real pose at each sampling period. Similarly, we calculate the average value of all absolute deviations between the reference heading angle and the robot heading angle for MAE. Lower error values correspond to better tracking accuracy. Additionally, we also compute the step-response characteristics for each of our designed controllers for more detailed analysis as shown in Fig. 9. On the other hand, we compare the tracking results for different prediction horizon values ($N_{p}$) in the case of Non-linear Model Predictive Control.
Table 2: Comparison of tracking error of omni drive robot for both type fuzzyPID and NMPC in absence of noise
T1-fuzzyPID IT2-fuzzyPID NPMC($N_{p}=15$)
Tracking time ME($XY$) MAE($\theta$) ME($XY$) MAE($\theta$) ME($XY$) MAE($\theta$)
20 seconds 0.0874 m 0.0888 rad 0.0855 m 0.0849 rad 0.0722 m 0.0625 rad
30 seconds 0.0521 m 0.05349 rad 0.0501 m 0.0535 rad 0.0439 m 0.0385 rad
6 Comparison of fuzzyPIDs and NMPC
6.1 In Absence of Noise
Table 2 presents the tracking error for both fuzzyPIDs and NMPC with a prediction horizon value of 15 in the absence of noise. The NMPC controller achieves a lower error compared to fuzzyPIDs in both metrics while a similar tracking performance is evaluated for fuzzy based controllers. As the tracking time increases, a corresponding improvement in tracking accuracy is observed for both controllers. On the other hand, we found fuzzy Controllers to be more computationally efficient than NMPC, albeit at the expense of tracking accuracy.
@@ 6.1 In Absence of Noise
Table 3: Comparison of tracking error of omni drive robot for both type fuzzyPIDs and NMPC in presence of noise
T1-fuzzyPID IT2-fuzzyPID NPMC($N_{p}=15$)
Tracking time ME($XY$) MAE($\theta$) ME($XY$) MAE($\theta$) ME($XY$) MAE($\theta$)
30 seconds 0.0657 m 0.0608 rad 0.0647 m 0.0588 rad 0.0566 m 0.0586 rad
6.2 Random Noise Injection
To assess the noise rejection capabilities of our proposed controllers for the Omni-drive robot, we introduced randomly generated noise into the feedback path of the input pose state at each timestamp, modeled by the equation $\dfrac{rand()}{6}\times\sin(\dfrac{nt_{s}}{5})$. Simulations were conducted over 30 seconds, with the resulting trajectory responses displayed in Fig. 12 and Fig. 12. The NMPC controller is found to be more robust to external noise compared to all other controllers. Additionally, the IT2-fuzzyPID controller exhibited greater noise rejection capability than the T1-fuzzyPID controller, attributable to its ability to model uncertainties, despite showing similar tracking performance in the absence of noise. The Mean Error (ME) and Mean Absolute Error (MAE) metrics for the proposed controllers in presence of noise are summarized in Table 3.
Metric T1-fuzzyPID IT2-fuzzyPID NMPC
$x$ $y$ $\theta$ $x$ $y$ $\theta$ $x$ $y$ $\theta$
$M_{\mathrm{p}}\%$ 10.800 9.448 78.904 10.611 9.692 77.887 5.083 3.480 95.766
$t_{r},0.1\rightarrow 0.9$ 0.991 1.186 0.215 0.949 1.133 0.242 2.252 0.914 0.175
$t_{s}\pm 10\%$ 1.610 1.186 2.282 1.522 1.133 2.160 0.921 0.914 1.514
Table 4: Step-response characteristics, such as overshoot($M_{\mathrm{p}}\%$), rise time($\boldsymbol{t}_{r}$) and settling time($\boldsymbol{t}_{s}$) of our proposed controllers
6.3 Step Response Analysis
The step-response analysis reveals that the NMPC controller achieves the fastest response, characterized by the minimal overshoot, shortest rise time and settling time, thereby ensuring superior stability. This performance is followed closely by the IT2-fuzzyPID controller, which outperforms the T1-fuzzyPID controller in trajectory tracking. Detailed insights into the step-response characteristics of all controllers are provided in Table 4.
Figure 11: Comparison of noise rejection capabilities of our controllers on pose(XY)
Figure 12: Comparison of noise rejection capabilities of our controllers on pose($\theta$)